Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
lumaca, nardecchia, berni - 11:20 Friday 12 April 2024 (63943) Print this report
WI CO2 chiller setpoint change

Following the activity of yesterday night (63940), we noticed that the WI main CO2 laser continued to be in an unstable state, so we decided to act on WI CO2 chiller setpoint, changing it from 19.02 °C to 19.04°C (08.27 UTC, Fig. 1). Also the BCK chiller set point has been changed at the same setpoint.

Soon after, we noticed a prominent change in the WI CO2 signals and we needed to recover the WI CO2 DAS outer ring power injected in the ITF, by rotating back the relative wave plate to recover the condition with respect to yesterday’s action (08.45 UTC, Fig. 2).

We’ll continue to monitor the situation to evaluate if a further chiller setpoint small change will be needed.

Images attached to this report
Search Help